OpenStudy (anonymous):

at which height from the earth urface at which acceleration due to gravity decreased by 75% of its value at earth surface?

OpenStudy (anonymous):

i tried it by using formula g=g/(1-h/R)^2 not getting my ans

OpenStudy (jamesj):

Well, in general F = ma and the force due to gravity acting on a mass m is F = - GMm/R^2 where G is the universal constant of gravity, M the mass of the earth and R the radius of the earth. Combining the two equations: for a mass m--like you or me or my cat--at the earth surface, it experiences an acceleration of a = - GM/R^2 Now ...

OpenStudy (jamesj):

The one thing we're varying is the radial distance from the center of the earth. Let's write our radial distance as r = R + h, where h is the height. Then the question is: For what value of h is \[ \frac{GM}{r^2} = \frac{3}{4} \frac{GM}{R^2} \]

OpenStudy (anonymous):

but it have no connection with this que we have to ue direct formula as in compettition only1 min is enough for one ques i think so..

OpenStudy (jamesj):

Well, this is how you solve it. Given that, what is an expression for h?

OpenStudy (anonymous):

for thi we had a formula but i m unable to solve this a it is becoming more complex gh=g/(1+h/r)^2

OpenStudy (jamesj):

I don't know where that formula comes from. Are you following my logic however?

OpenStudy (anonymous):

acc. to ur solution we cant find the heigh as we have totake out an in a comparision way am i rit sir?? actually i m v.bad in physics frankly sayin

OpenStudy (jamesj):

We can find it. r = R + h. Hence, from using the last equation I wrote down, \[ r^2 = \frac{4}{3}R^2 \] \[ (R+h)^2 = \frac{4}{3}R^2 \] \[ R + h = \sqrt{4/3} R \] Now solve that for h. I strongly recommend you read what I wrote above and ask me questions about it. What I have written down is exactly the logic for getting to the answer.

OpenStudy (anonymous):

i tried sir but not getting ans by dis too :( the ans is h=Re Re is radius of earth surface

OpenStudy (jamesj):

Do you agree F = ma?

OpenStudy (anonymous):

yes and what ur saying i already did that we had a derivetion by which concluded the formula to find h at altitude that is gh=g/(1+h/r)^2 do u want to see the derivation??

OpenStudy (jamesj):

Where is 3/4 or 75% in your formula?

OpenStudy (anonymous):

plz solve this que by your method firt if i had objection i ll ask u ok

OpenStudy (anonymous):

its a general formula sir by which we can use it every where

OpenStudy (jamesj):

I have solved it for you. There's one small step of algebra left. Read my solution and ask me questions about it if you like.

OpenStudy (anonymous):

but the ans is Re and what the way u said ans is not coming :(

OpenStudy (jamesj):

Oh, misread the question. The acceleration is decreased BY 75%, not to 75%. My mistake. If it is decreased by 75%, then it 1/4 of what it was, hence \[ \frac{GM}{r^2} = \frac{1}{4} \frac{GM}{R^2} \] which implies \[ r^2 = 4 R^2 \] i.e., \[ (R+h)^2 = 4 R^2 \] \[ R + h = 2R \] \[ h = R \]

OpenStudy (anonymous):

thanQ..

OpenStudy (anonymous):

had one more ques hold on i m posting it

OpenStudy (jamesj):

post it on the left